OpenStudy (anonymous):

Marcy and Lucia had lunch at a restaurant. The bill for the lunch was $24 and they gave the waitress an 18% tip. Whats the total cost of their lunch?

OpenStudy (texaschic101):

24 + 24(.18) = 24 + 4.32 = 28.32

OpenStudy (anonymous):

Thank you, do you think you can help me with this last question?

OpenStudy (anonymous):

A widescreen television sells for $1050. The sales tax is 8%. What is the total cost of the television?

OpenStudy (texaschic101):

you mean the last one you posted ?

OpenStudy (texaschic101):

1050 + 1050(.08) = 1050 + 84 = $ 1134

OpenStudy (anonymous):

Thank you :)

OpenStudy (texaschic101):

sure thing :)